# Code: https://github.com/icey-zhang/EfficientMFD

### v1

@@ 5 Conclusion
6 Appendix
6.1 More Fusion Visualization Results
@@
More comparisons of infrared-visible image fusion visualization results are depicted in Figure 6, 7 and 8. These fusion results demonstrate the advantages of synchronously optimizing fusion and detection tasks. With minimal training costs, we obtain fusion images that are visually and detection-friendly. Specifically, the fusion images retain significant object information extracted from infrared images, while also preserving detailed information such as texture, color, and background from visible images. Our method effectively combines the strengths of both modalities to enhance the overall performance of the detection.
Figure 6: Visualize results of image fusion on M3FD.
Figure 7: Visualize results of image fusion on RoadScene.
Figure 8: Visualize results of image fusion on TNO.
6.2 More Detection Visualization Results
@@
The visualization of infrared-visible object detection are demonstrated in Figure 9. Our fusion images are noticeably superior due to the effective combination of object information from infrared images and texture details from visible images. This integration enables the object detection network to distinguish between the background and the object clearly. Additionally, for small and occluded objects, the clear edge details assist the network in achieving improved detection results.
Figure 9: Visualize results of object detection on M3FD.

### v2

@@ 5 Conclusion
Appendix A Appendix
A.1 Metric
Three metrics are used for MF evaluation: entropy (EN) Roberts et al. [2008], mutual information (MI) Qu et al. [2002] and visual information fidelity (VIF) Han et al. [2013]. EN evaluates the information richness in an image, and the higher EN means more information. MI evaluates the information similarity between the input images and fused images. The higher MI illustrates more information of the input images is fused. VIF measures the ability to extract visible information from the input image, and a larger VIF represents less visible distortion in the fused result. Here, we use the V channel in the HSV space of the fusion results to calculate these metrics. Moreover, we use $\text{mAP}_{50:95}$ He and Todorovic [2022] to comprehensively evaluate OD performance, where the average of mAPs sampling every 5 from AP50 to AP95 is calculated. A higher $\text{mAP}_{50:95}$ means a better OD effect.
A.2 Visialization of Feature Maps
We have implemented the Object-Region-Pixel Phylogenetic Tree (ORPPT) to explore the hierarchical interactions under different granularity views and extract various features across multiple region scales, as introduced in Section 3.2. As shown in Figure 6, the detailed information will decrease as the backbone structure shared by the detection network deepens, resulting in a decrease in detailed information. This may affect the absorption of these features by the fusion network of the shared backbone network, thereby affecting pixel-level fusion tasks. This provides evidence for our analysis of the reasons for Section 5 ablation experiment for Object-Region-Pixel Phylogenetic Tree. This illustrates the importance of balancing the semantic information provided between the OD and MF network with pixel-level information.
Figure 6: Feature map visualization of various branches in the ORPPT.
A.3 More Ablation Studies
To compare the effectiveness of different Multi-task learning (MTL) methods on our algorithm, we selected three robust MTL optimization methods. As shown in Table 6, all MTL methods addressed the conflict between the MF and OD tasks to varying degrees. By introducing the concept of Gradient Matrix Task-Alignment (GMTA), we achieved a better balance in the gradient optimization process between the two tasks, resulting in the best performance across all metrics.
Table 6: Ablation of different MTL methods on M3FD.
Method EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
E2E-MFD (w/o GMTA) 6.12 14.70 1.39 90.05 62.60
PCGradYu et al. [2020] 6.13 15.01 1.48 90.59 62.71
CAGradLiu et al. [2021c] 6.17 15.05 1.48 90.71 62.74
Nash-MTLNavon et al. [2022] 6.29 15.28 1.51 90.91 62.97
E2E-MFD (w GMTA) 6.36 15.47 1.65 91.80 63.83
A.4 Experiments on DroneVehicle
DroneVehicle comprises aerial RGB-IR images captured by drones, encompassing various scenes from an aerial perspective. It contains five categories of target objects. The dataset consists of 28,439 pairs of images, divided into training, validation, and test sets. We train the MF network on the training set and perform inference directly on the test set. Finally, fusion images are generated for both the training and test sets. All detection accuracies are obtained on the test set. The pretrained LSKNet is used for the initialization of the object detection network, and we fine-tune it for 12 epochs with a batch size of 4. To fully demonstrate the effectiveness of our method, we compare some state-of-the-art (SOTA) models, and experiments are also conducted on YOLOv5s-OBB using the generated fusion images, all images are resized to 640$\times$640 and trained from scratch for 50 epochs with a batch size of 64.
The visualization results of the fusion and detection of DroneVehicle are shown in Figure 7 and 8. In the process of generating fusion images, we first obtain RGB images, then convert them to HSV format, and only take the V channel as the final fusion image. It preserves the brightness information of the objects while containing rich details from visible images. As shown in Figure 7, the first row and the second row are images extracted from the training set and the test set, respectively. Although the style of the fusion images resembles that of the infrared images, by combining visible image information, the fusion images can differentiate different parts of each object, reflecting varying degrees of brightness information. This is advantageous for fine-grained classification. Additionally, as shown in Figure 8, the visually friendly fusion images also assist in the object detection task, allowing for sufficient detection even in cases of small and dense objects.
Figure 7: Qualitative results of image fusion on DroneVehicle.
Figure 8: Visual results of object detection on DroneVehicle.
A.5 More Fusion Visualization Results
@@
More comparisons of infrared-visible image fusion visualization results are depicted in Figure 9, 10 and 11. These fusion results demonstrate the advantages of synchronously optimizing fusion and detection tasks. With minimal training costs, we obtain fusion images that are visually and detection-friendly. Specifically, the fusion images retain significant object information extracted from infrared images, while also preserving detailed information such as texture, color, and background from visible images. Our method effectively combines the strengths of both modalities to enhance the overall performance of the detection.
Figure 9: Qualitative results of image fusion on M3FD.
Figure 10: Qualitative results of image fusion on RoadScene.
Figure 11: Qualitative results of image fusion on TNO.
A.6 More Detection Visualization Results
@@
The visualization of infrared-visible object detection is demonstrated in Figure 12. Our fusion images are noticeably superior due to the effective combination of object information from infrared images and texture details from visible images. This integration enables the object detection network to distinguish between the background and the object clearly. Additionally, for small and occluded objects, the clear edge details assist the network in achieving improved detection results.
Figure 12: Visual results of object detection on M3FD.
A.7 Limitations and Broader Impacts
Limitations. Our E2E-MFD approach is effective for the joint learning of the multimodal fusion and object detection task and has been validated on various datasets. However, the validations of the current model rely on the visible and infrared modalities. Constrained by the scarcity of relevant datasets within the community, the paper lacks validation with additional datasets containing new modalities. Future research will focus on addressing this gap by exploring, constructing, and incorporating more diverse multimodal datasets serving multimodal fusion detection.
Broader Impacts. Our paper aims to broaden the applicability of joint learning of multimodal data and object detection to various research domains. However, this broader scope may present challenges when using the model in domains that include harmful content. These challenges arise from the data itself, rather than from the model. Therefore, it is crucial to have adequate data regularization to effectively address these concerns.